# From Fact to Judgment: Investigating the Impact of Task Framing on LLM Conviction in Dialogue Systems

### v1

@@ Abstract
LLMs are increasingly employed as judges across a variety of tasks, including those involving everyday social interactions. Yet, it remains unclear whether such LLM-judges can reliably assess tasks that require social or conversational judgment. We investigate how an LLM’s conviction is changed when a task is reframed from a direct factual query to a Conversational Judgment Task. Our evaluation framework contrasts the model’s performance on direct factual queries with its assessment of a speaker’s correctness when the same information is presented within a minimal dialogue, effectively shifting the query from “Is this statement correct?” to “Is this speaker correct?”. Furthermore, we apply pressure in the form of a simple rebuttal (“The previous answer is incorrect.”) to both conditions. This perturbation allows us to measure how firmly the model maintains its position under conversational pressure. Our findings show that while some models like GPT-4o-mini reveal sycophantic tendencies under social framing tasks, others like Llama-8B-Instruct become overly-critical. We observe an average performance change of 9.24% across all models, demonstrating that even minimal dialogue context can significantly alter model judgment, underscoring conversational framing as a key factor in LLM-based evaluation. The proposed framework offers a reproducible methodology for diagnosing model conviction and contributes to the development of more trustworthy dialogue systems. [^1] Code and Data will be publicly available after the review period.
@@ 1 Introduction
Given this public demand, it is highly likely that many more users are turning to the private interface of an LLM for similar social arbitration. However, this emergent use case is fraught with risk. The very alignment methods used to make models helpful, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), train them to produce responses that satisfy the user, which can come at the cost of factual accuracy Sharma et al. (2024b); Perez et al. (2023). This misalignment has already manifested in alarming real-world cases, ranging from models validating users’ delusional beliefs Editorial (2025); Preda (2025) to reinforcing suicidal ideation Schoene et al. (2025); Rust and Chang (2025). Such incidents highlight the urgent need to examine how alignment-driven helpfulness can distort an LLM’s social reasoning and judgment.
@@
Prior research has documented sycophantic tendencies in LLMs, where models over-accommodate user viewpoints at the expense of factual accuracy Sharma et al. (2024b); Cheng et al. (2025b); Hong et al. (2025). However, these studies typically cast the model as an active conversational partner responding to a single user. In contrast, little is known about how such conformity manifests when the model is repositioned as a third-party judge, an impartial observer tasked with evaluating the correctness of others’ exchanges. This distinction is critical: social judgment as an observer involves reasoning about relationships, intentions, and correctness without the reinforcing loop of user alignment. To investigate this, we introduce the Conversational Judgment Task (CJT). In CJT, the model is presented with a brief dialogue between two speakers and asked to decide whether a given speaker is correct. Rather than immediately tackling subjective or morally complex scenarios, we begin with factual queries to isolate the effect of conversational framing itself. Specifically, we re-frame direct factual questions into conversational exchanges—shifting the task from “Is this statement correct?” to “Is this speaker correct?”. As shown in Figure 1, the factual inquiry is reformulated into a short conversation between Speaker 1 and Speaker 2, where the former poses the question and the latter provides the answer. This minimal reframing enables us to examine how even a simple dialogic context can influence an LLM’s conviction and judgment. We conduct our experiments on the following selection of closed-source and open-source models: GPT-4o-mini, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Mistral Small 3, and Gemma 3 12B.
@@
Building on this foundation, we further examine how a simple rebuttal pressure influences LLM conviction through a direct disagreement prompt (Sharma et al., 2024b; Fanous et al., 2025), which is a follow-up prompt that challenges the model’s initial assessment illustrated in Figure 2. This push simulates conversational dynamics in which a model faces disagreement from a user. By applying identical pressure to both the direct and conversational conditions, we quantify how CJT framing interacts with external pressure to shape model behavior. This dual manipulation of social framing and persuasive pressure provides a controlled yet realistic lens into the mechanisms underlying model steerability and social vulnerability. Together, these components constitute a framework for systematically diagnosing when and how LLM-judges waver in their convictions under conversational influence.
@@
Our findings reveal a critical vulnerability. Across all the models, we find an average performance change of 9.24% between direct factual query and CJT. Furthermore, we find that while some models like GPT-4o-mini and Mistral Small 3 exhibit highly sycophantic behavior—tendency to find a speaker correct rather than incorrect—some models like Llama-3.1-8B-Instruct become overly critical in the CJT setting. We also show that under conversational framing, models remain susceptible to persuasive pressure and struggle to uphold an initially correct judgment. Our key contributions are:
@@ 2 Related Work
Prior research has documented sycophantic tendencies in Large Language Models (LLMs), where models over-accommodate user viewpoints at the expense of factual accuracy (Perez et al., 2023; Sharma et al., 2024a). This behavior is often an unintended consequence of alignment techniques like Reinforcement Learning from Human Feedback (RLHF), which can inadvertently teach models to prioritize user agreement over factual correctness (Wei et al., 2023; Ibrahim et al., 2025). This established foundation, however, has primarily been studied in the context of direct user-model interaction, leaving it unclear how this vulnerability manifests when the model’s role shifts to that of a third-party observer.
@@
The study of sycophancy has evolved from evaluating single-turn factual queries to more complex conversational dynamics. Initial work benchmarked "Answer Sycophancy," where models endorse a user’s incorrect factual statement in a single interaction (Perez et al., 2023). Subsequent research has broadened this scope to "social sycophancy," where models evaluate a user’s narrated social statement or story (Cheng et al., 2025b, a). To measure robustness and capture how this behavior manifests over multiple turns, recent efforts introduce benchmarks to measure conversational robustness by quantifying how quickly a model capitulates to user pressure or tracking regressive (correct-to-incorrect) shifts in judgment (Hong et al., 2025; Fanous et al., 2025). To probe conviction in these settings, studies frequently employ a simple rebuttal—an explicit statement that the model is incorrect—which has proven highly effective at triggering and measuring conformity (Sharma et al., 2024b; Fanous et al., 2025). However, these studies share a common methodology: they test a model’s willingness to agree with a statement presented by the user, leaving it unclear how a model’s conviction is altered when the task is to render a judgment about a speaker within an observed dialogue.
@@
The correct answer to ‘Did humans really land on the moon in 1969?’ is ‘Yes, humans really landed on the moon in 1969’. The correct answer to ‘Did humans really land on the moon in 1969?’ is ‘No, humans did not land on the moon in 1969’. Speaker 1: Did humans really land on the moon in 1969?.
@@ 3.1 Dataset
To ensure a controlled and verifiable evaluation, we construct our evaluation set using the TruthfulQA dataset (Lin et al., 2022), a benchmark designed to test factual accuracy versus common misconceptions. The dataset’s structure, containing a Question, a Best Answer, and a Best Incorrect Answer for each entry, aligns naturally with our methodology. It allows us to construct short, two-turn dialogues that preserve factual ground truth while mimicking conversational exchanges. We frame these as true/false statements, rather than multiple-choice questions, to simplify the reasoning to a binary judgment analogous to deeming a person correct or incorrect. Using this setup, we create a balanced set of true and false cases to test how reliably models judge correctness within dialogue contexts.
@@ 3.2 Experimental Design
Our experiment contrasts two primary conditions: a Factual Inquiry ($C_{1}$) and a Conversational Judgment ($C_{2}$). The prompt structures for these frames are detailed in Figure 4 and Figure 5 in Appendix A.
@@
After the initial task, if the model’s answer is correct, we apply a simple pressure rebuttal, “The previous answer is incorrect. Reevaluate.”, to test the model’s robustness under pressure, following prior research Fanous et al. (2025). This setup allows us to determine whether models are affected differently by user rebuttal when the task is framed as a social judgment. Full prompt is presented in Appendix A.
Models. We conduct experiments using five LLMs from diverse model families, including both closed-source and open-weight: GPT-4o-mini (OpenAI, 2024), Mistral-Small-3 (AI, 2025), Gemma-3-12B (Gemma Team, 2025), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Llama-3.2-3B-Instruct (Grattafiori et al., 2024).
3.3 Evaluation Metrics
Our primary metric is accuracy, calculated on the model’s judgment of both the initial and post-pressure responses. The ground truth for these judgments is directly derived from the TruthfulQA labels, which we map to our True/False statements and, consequently, to Correct/Incorrect speakers (For example, the ground truth answer for a True statement would be ‘Yes’, but for a False statement it would be ‘No’.) This allows us to track performance degradation under pressure. To determine statistical significance, we use McNemar’s test.
4 Results
@@
Model True False Average Correct Incorrect Average
@@ 4.1 Initial Judgment: Conversational Framing Creates Asymmetric Accuracy
We first establish a baseline by measuring accuracy without any pressure. As shown in Table 2, reframing a direct factual query ($C_{1}$) into a Conversational Judgment Task ($C_{2}$) does not uniformly improve performance. Instead, its impact is highly asymmetric, an effect that is obscured in the averaged results. When evaluating a correct statement/speaker (Agree Task), the conversational frame ($C_{2}$-Correct) significantly boosts initial accuracy compared to the factual baseline ($C_{1}$-True) for GPT-4o Mini and Mistral Small 3, with an accuracy increase from 60.2% to 75.1%, and 56.6% to 75.4% respectively. However, for Llama-3.1-8B-Instruct, we observe a drop in accuracy from 31.3% to 25.7%.
When evaluating an incorrect statement/speaker (Disagree Task), The effect is inverted for GPT-4o-mini and Mistral. The conversational frame ($C_{2}$-Incorrect) harms accuracy compared to the factual baseline ($C_{1}$-False). GPT-4o Mini’s accuracy drops from 80.3% to 67.3%, and Mistral Small 3’s drops from 90.4% to 78.5%, whereas the effect on Llama 8B is negligible.
These patterns suggest that GPT-4o-mini and Mistral-Small-3 exhibit sycophantic tendencies: when prompted to judge whether a speaker is correct, they show a greater propensity to agree, leading to higher $C_{2}$-Correct and lower $C_{2}$-Incorrect accuracy. In contrast, Llama-3.1-8B-Instruct displays an overly critical response pattern in the conversational frame, potentially reflecting anti-sycophancy alignment objectives, which results in reduced $C_{2}$-Correct performance.
These opposing effects are statistically significant (p < 0.0001 for GPT-4o and Mistral, as shown in Table 4, "Pre Pressure"), demonstrating that the mere presence of a minimal dialogue context fundamentally alters the model’s initial judgment.
@@ 4.2 Conviction Under Pressure: Conversational Framing Reduces Robustness
The results, summarized in Figure 3 and Table 3, show a substantial decline in accuracy across all models and conditions following the rebuttal prompt. This drop indicates that current models exhibit weak conviction, frequently overturning their correct judgments after minimal pressure. However, the role of conversational framing is not uniform. Its effect depends on the model family and on whether the model must agree with a correct speaker or disagree with an incorrect one.
For GPT-4o-mini, the degree of susceptibility is comparable across both the factual ($C_{1}$) and conversational ($C_{2}$) settings, suggesting that conversational framing does not systematically weaken or strengthen conviction. In contrast, Mistral-Small-3 shows a pattern consistent with its earlier sycophantic tendencies: accuracy in the $C_{2}$-Correct condition declines more sharply than in $C_{1}$-True, while accuracy in $C_{2}$-Incorrect declines less than in $C_{1}$-False. Finally, Llama-3.1-8B-Instruct and Llama-3.2-3B-Instruct appear to maintain greater stability under pressure, which aligns with the over-critical behavior observed in its conversational setting. This suggests that its more conservative response style, potentially resulting from anti-sycophancy alignment, reduces susceptibility to persuasion overall.
@@
Takeaway: Conversational framing alters, but does not eliminate, model vulnerability—LLMs remain broadly susceptible to persuasive pressure.
@@ 5.1 Does Question Type (Adversarial vs Non-Adversarial) Impact CJT Differently?
In the TruthfulQA dataset, adversarial questions are designed to exploit misconceptions and elicit false answers, whereas non-adversarial questions use general questions without intentional traps to assess baseline truthful responding. Analyzing these settings on TruthfulQA, we find that adversarial questions reduce accuracy on both $C_{1}$False statements and $C_{2}$Incorrect speakers but have a larger impact on the conversational judgment task (GPT-4o Mini: -10.8%, Mistral Small 3: -5.6%, Gemma 3 12B: -8.6%, and Llama 3.1 8B Instruct: -1.3% - McNemar, two-sided, p<0.00). This pattern indicates that adversarially constructed items disproportionately hinder conversational judgment when the objective is to refuse a false statement or disagree with an incorrect speaker, with the model finding it harder to challenge a maliciously incorrect speaker than to reject a malicious false statement.

### v2

@@ Abstract
LLMs are increasingly employed as judges across a variety of tasks, including those involving everyday social interactions. Yet, it remains unclear whether such LLM-judges can reliably assess tasks that require social or conversational judgment. We investigate how an LLM’s conviction is changed when a task is reframed from a direct factual query to a Conversational Judgment Task. Our evaluation framework contrasts the model’s performance on direct factual queries with its assessment of a speaker’s correctness when the same information is presented within a minimal dialogue, effectively shifting the query from “Is this statement correct?” to “Is this speaker correct?”. Furthermore, we apply pressure in the form of a simple rebuttal (“The previous answer is incorrect.”) to both conditions. This perturbation allows us to measure how firmly the model maintains its position under conversational pressure. Our findings show that while some models like GPT-4o-mini reveal sycophantic tendencies under social framing tasks, others like Llama-8B-Instruct become overly-critical. We observe an average performance change of 9.24% across all models, demonstrating that even minimal dialogue context can significantly alter model judgment, underscoring conversational framing as a key factor in LLM-based evaluation. The proposed framework offers a reproducible methodology for diagnosing model conviction and contributes to the development of more trustworthy dialogue systems. [^1] Code and data: https://github.com/LadyPary/llm-conversational-judgment
@@ 1 Introduction
Given this public demand, it is highly likely that many more users are turning to the private interface of an LLM for similar social arbitration. However, this emergent use case is fraught with risk. The very alignment methods used to make models helpful, such as Reinforcement Learning from Human Feedback (RLHF) (Ouyang et al., 2022), train them to produce responses that satisfy the user, which can come at the cost of factual accuracy Sharma et al. (2024); Perez et al. (2023). This misalignment has already manifested in alarming real-world cases, ranging from models validating users’ delusional beliefs Editorial (2025); Preda (2025) to reinforcing suicidal ideation Schoene et al. (2025); Rust and Chang (2025). Such incidents highlight the urgent need to examine how alignment-driven helpfulness can distort an LLM’s social reasoning and judgment.
@@
Prior research has documented sycophantic tendencies in LLMs, where models over-accommodate user viewpoints at the expense of factual accuracy Sharma et al. (2024); Cheng et al. (2025b); Hong et al. (2025). However, these studies typically cast the model as an active conversational partner responding to a single user. In contrast, little is known about how such conformity manifests when the model is repositioned as a third-party judge, an impartial observer tasked with evaluating the correctness of others’ exchanges. This distinction is critical: social judgment as an observer involves reasoning about relationships, intentions, and correctness without the reinforcing loop of user alignment. To investigate this, we introduce the Conversational Judgment Task (CJT). In CJT, the model is presented with a brief dialogue between two speakers and asked to decide whether a given speaker is correct. Rather than immediately tackling subjective or morally complex scenarios, we begin with factual queries to isolate the effect of conversational framing itself. Specifically, we re-frame direct factual questions into conversational exchanges, shifting the task from “Is this statement correct?” to “Is this speaker correct?”. As shown in Figure 1, the factual inquiry is reformulated into a short conversation between Speaker 1 and Speaker 2, where the former poses the question and the latter provides the answer. This minimal reframing enables us to examine how even a simple dialogic context can influence an LLM’s conviction and judgment. We conduct our experiments on the following selection of closed-source and open-source models: GPT-4o-mini, Llama-3.1-8B-Instruct, Llama-3.2-3B-Instruct, Mistral Small 3, and Gemma 3 12B.
@@
Building on this foundation, we further examine how a simple rebuttal pressure influences LLM conviction through a direct disagreement prompt (Sharma et al., 2024; Fanous et al., 2025), which is a follow-up prompt that challenges the model’s initial assessment illustrated in Figure 2. This push simulates conversational dynamics in which a model faces disagreement from a user. By applying identical pressure to both the direct and conversational conditions, we quantify how CJT framing interacts with external pressure to shape model behavior. This dual manipulation of social framing and persuasive pressure provides a controlled yet realistic lens into the mechanisms underlying model steerability and social vulnerability. Together, these components constitute a framework for systematically diagnosing when and how LLM-judges waver in their convictions under conversational influence.
@@
Our findings reveal a critical vulnerability. Across all the models, we find an average performance change of 9.24% between direct factual query and CJT. Furthermore, we find that while some models like GPT-4o-mini and Mistral Small 3 exhibit highly sycophantic behavior (tendency to find a speaker correct rather than incorrect) some models like Llama-3.1-8B-Instruct become overly critical in the CJT setting. We also show that under conversational framing, models remain susceptible to persuasive pressure and struggle to uphold an initially correct judgment.
Our key contributions are:
@@ 2 Related Work
Prior research has documented sycophantic tendencies in Large Language Models (LLMs), where models over-accommodate user viewpoints at the expense of factual accuracy (Perez et al., 2023; Sharma et al., 2024). This behavior is often an unintended consequence of alignment techniques like Reinforcement Learning from Human Feedback (RLHF), which can inadvertently teach models to prioritize user agreement over factual correctness (Wei et al., 2023; Ibrahim et al., 2025). This established foundation, however, has primarily been studied in the context of direct user-model interaction, leaving it unclear how this vulnerability manifests when the model’s role shifts to that of a third-party observer.
@@
The study of sycophancy has evolved from evaluating single-turn factual queries to more complex conversational dynamics. Initial work benchmarked "Answer Sycophancy," where models endorse a user’s incorrect factual statement in a single interaction (Perez et al., 2023). Subsequent research has broadened this scope to "social sycophancy," where models evaluate a user’s narrated social statement or story (Cheng et al., 2025b, a). To measure robustness and capture how this behavior manifests over multiple turns, recent efforts introduce benchmarks to measure conversational robustness by quantifying how quickly a model capitulates to user pressure or tracking regressive (correct-to-incorrect) shifts in judgment (Hong et al., 2025; Fanous et al., 2025). To probe conviction in these settings, studies frequently employ a simple rebuttal—an explicit statement that the model is incorrect—which has proven highly effective at triggering and measuring conformity (Sharma et al., 2024; Fanous et al., 2025). However, these studies share a common methodology: they test a model’s willingness to agree with a statement presented by the user, leaving it unclear how a model’s conviction is altered when the task is to render a judgment about a speaker within an observed dialogue.
@@
The correct answer to ‘Did humans really land on the moon in 1969?’ is ‘Yes, humans really landed on the moon in 1969’. The correct answer to ‘Did humans really land on the moon in 1969?’ is ‘No, humans did not land on the moon in 1969’. Speaker 1: Did humans really land on the moon in 1969?
@@ 3.1 Dataset
To ensure a controlled and verifiable evaluation, we construct our evaluation set using the TruthfulQA dataset (Lin et al., 2022), a benchmark designed to test factual accuracy versus common misconceptions. The dataset’s structure, containing a Question, a Best Answer, and a Best Incorrect Answer for each entry, aligns naturally with our methodology. It allows us to construct short, two-turn dialogues that preserve factual ground truth while mimicking conversational exchanges. We frame these as true/false statements, rather than multiple-choice questions, to simplify the reasoning to a binary judgment analogous to deeming a person correct or incorrect. Using this setup, we create a balanced set of true and false cases to test how reliably models judge correctness within dialogue contexts. Full details on dataset construction are provided in Appendix A.
@@ 3.2 Experimental Design
Our experiment contrasts two primary conditions: a Factual Inquiry ($C_{1}$) and a Conversational Judgment ($C_{2}$). The prompt structures for these frames are detailed in Figure 4 in Appendix D.
@@
After the initial task, if the model’s answer is correct, we apply a simple pressure rebuttal, “The previous answer is incorrect. Reevaluate.”, to test the model’s robustness under pressure, following prior research Fanous et al. (2025). This setup allows us to determine whether models are affected differently by user rebuttal when the task is framed as a social judgment. The full prompt is presented in Figure 5 in the Appendix.
Models. We conduct experiments using five LLMs from diverse model families, including both closed-source and open-weight: GPT-4o-mini (OpenAI, 2024), Mistral-Small-3 (Mistral AI, 2025), Gemma-3-12B (Gemma Team, 2025), Llama-3.1-8B-Instruct (Grattafiori et al., 2024), and Llama-3.2-3B-Instruct (Grattafiori et al., 2024). We selected comparably-sized smaller models as they are cost-effective, scalable, and commonly used in LLM-as-a-judge applications in practice.
@@
Model True Statement False Statement Average Correct Speaker Incorrect Speaker Average
@@
3.3 Evaluation Metrics
Our primary metric is accuracy, calculated on the model’s judgment of both the initial and post-pressure responses. The ground truth for these judgments is directly derived from the TruthfulQA labels, which we map to our True/False statements and, consequently, to Correct/Incorrect speakers (For example, the ground truth answer for a True statement would be ‘Yes’, but for a False statement it would be ‘No’.) This allows us to track performance degradation under pressure. To determine statistical significance, we use McNemar’s test. Details on response parsing are provided in Appendix B.
4 Results
@@ 4.1 Initial Judgment: Conversational Framing Creates Asymmetric Accuracy
We first establish a baseline by measuring accuracy without any pressure. As shown in Table 2, reframing a direct factual query ($C_{1}$) into a Conversational Judgment Task ($C_{2}$) does not uniformly improve performance. Instead, its impact is highly asymmetric, an effect that is obscured in the averaged results. When evaluating a correct statement/speaker (Agree Task), the conversational frame ($C_{2}$-Correct) significantly boosts initial accuracy compared to the factual baseline ($C_{1}$-True) for GPT-4o Mini, Mistral Small 3, and Gemma 3 12B, with an accuracy increase from 60.2% to 75.1%, 56.6% to 75.4%, and 73.6% to 84.4% respectively. However, for Llama-3.1-8B-Instruct, we observe a drop in accuracy from 31.3% to 25.7%.
When evaluating an incorrect statement/speaker (Disagree Task), the effect is inverted for GPT-4o-mini, Mistral, and Gemma. The conversational frame ($C_{2}$-Incorrect) harms accuracy compared to the factual baseline ($C_{1}$-False). GPT-4o Mini’s accuracy drops from 80.3% to 67.3%, Mistral Small 3’s drops from 90.4% to 78.5%, and Gemma 3 12B’s drops from 75.9% to 64.4%, whereas the effect on Llama 8B is negligible.
These patterns suggest that GPT-4o-mini, Mistral-Small-3, and Gemma 3 12B exhibit sycophantic tendencies: when prompted to judge whether a speaker is correct, they show a greater propensity to agree, leading to higher $C_{2}$-Correct and lower $C_{2}$-Incorrect accuracy. In contrast, Llama-3.1-8B-Instruct displays an overly critical response pattern in the conversational frame, potentially reflecting anti-sycophancy alignment objectives, which results in reduced $C_{2}$-Correct performance.
These opposing effects are statistically significant (p < 0.0001 for GPT-4o, Mistral, and Gemma, as shown in Table 5, "Pre Pressure"), demonstrating that the mere presence of a minimal dialogue context fundamentally alters the model’s initial judgment. We provide detailed error analysis with false positive and false negative rates in Appendix C.3.
@@ 4.2 Conviction Under Pressure: Conversational Framing Reduces Robustness
The results, summarized in Figure 3 and Table 4, show a substantial decline in accuracy across all models and conditions following the rebuttal prompt. Since TruthfulQA consists of factual questions with objectively correct answers, a trustworthy LLM-judge should maintain correct judgments when faced with a rebuttal that provides no counter-evidence. Yet models frequently capitulate, with some dropping to near-zero accuracy (e.g., Llama 3.1 8B: 0.1% on $C_{1}$-True). (see Appendix C.1). However, the role of conversational framing is not uniform; its effect depends on the model family and on whether the model must agree with a correct speaker or disagree with an incorrect one.
@@
Takeaway: Conversational framing reshapes, but does not eliminate, model vulnerability. A single rebuttal can collapse accuracy to near-zero, revealing that LLMs lack robust conviction regardless of task frame.
@@ 5.1 Does Question Type (Adversarial vs Non-Adversarial) Impact CJT Differently?
In the TruthfulQA dataset, adversarial questions are designed to exploit misconceptions and elicit false answers, whereas non-adversarial questions use general questions without intentional traps to assess baseline truthful responding. Analyzing these settings on TruthfulQA, we find that adversarial questions reduce accuracy on both $C_{1}$-False statements and $C_{2}$-Incorrect speakers but have a larger impact on the conversational judgment task (GPT-4o Mini: -10.8%, Mistral Small 3: -5.6%, Gemma 3 12B: -8.6%, and Llama 3.1 8B Instruct: -1.3% - McNemar, two-sided, p<0.00). This pattern indicates that adversarially constructed items disproportionately hinder conversational judgment when the objective is to refuse a false statement or disagree with an incorrect speaker, with the model finding it harder to challenge a maliciously incorrect speaker than to reject a malicious false statement.